OpenStudy (anonymous):

what is the least common multiple of 12 and 18

OpenStudy (austinl):

\(12 = 2 \times 2\times 3\) \(18 = 2 \times 3 \times 3\) --------------------- The LCM of 12 , 18 is: \(2 \times 2 \times 3 \times 3=\) \(= 36\)